OpenStudy (rogue):

Can someone walk me through on solving this DE? The general solution to the differential equation y' = y + x^2 is y =

OpenStudy (amistre64):

get the y parts to one side

OpenStudy (bahrom7893):

integrating factors?

OpenStudy (bahrom7893):

y' - x^2 = y

OpenStudy (amistre64):

y' -y = x^2 then solve the characteristic eq r-1 = 0 when r = 1 \[y=c_1e^x+yp\] maybe

OpenStudy (amistre64):

or since its a first order we can go to the reversal of int by parts

OpenStudy (rogue):

Ahh, nevermind, I realize my mistake. Thanks anyways =)

OpenStudy (bahrom7893):

wait i moved the wrong thing over lol y' - y = x^2 e^(Int(-1)) = e^(-x) mehh lol im too slow haha

OpenStudy (amistre64):

\[y' -y = x^2\] \[e^{-ln(x)}(y' -y = x^2)\] \[e^{-ln(x)}y =\int x\ dx\] \[y/x =\frac{1}{2}x^2+c\] \[y =\frac{1}{2}x^3+Cx\] might be one solution

OpenStudy (amistre64):

the wolf like the first try

OpenStudy (rogue):

I got this \[y' - y = x^2\]\[y' e^{-x} + ye^{-x} = x^2 e^{-x} \rightarrow \frac {d}{dx} \left[ ye^{-x} \right] = x^2 e^{-x}\]\[\int\limits \frac {d}{dx} \left[ ye^{-x} \right] dx = \int\limits x^2 e^{-x} dx\]\[y e^{-x} = -e^{-x} (x^2 + 2x +2) + C\]\[y = Ce^{x} -x^2 - 2x - 2 \]

OpenStudy (amistre64):

thats it

OpenStudy (amistre64):

lol, I did something odd; shoulda been \(\int -1 dx=-x\)

OpenStudy (amistre64):

\[yp = ax^2+bx+c\] \[yp' = 2ax+b\] \[yp'-yp = 2ax+b-ax^2-bx-c=x^2\] \[\begin{matrix} (-a)x^2 & (2a-b)x&(b-c)\\ (1)x^2&0(x)&0\\ a=-1&b=-2&c=-2 \end{matrix}\] \[yp = -x^2-2x-2\] \[y=c_1e^x-x^2-2x-2 \]

OpenStudy (rogue):

Oh, thats interesting! thank you =)